### v3

# Collaborating with AI Agents: A Field Experiment on Teamwork, Productivity, and Performance

Abstract
We examined the mechanisms underlying productivity and performance gains from AI agents using a large-scale experiment on Pairit, a platform we developed to study human-AI collaboration. We randomly assigned 2,234 participants to human-human and human-AI teams that produced 11,024 ads for a think tank. We evaluated the ads using independent human ratings and a field experiment on X which garnered $\sim$5M impressions. We found human-AI teams produced 50% more ads per worker and higher text quality, while human-human teams produced higher image quality, suggesting a jagged frontier of AI agent capability. Human-AI teams also produced more homogeneous outputs (or diversity collapse) as their ads were more self-similar. The field experiment revealed higher text quality (produced by human-AI teams) improved click-through rates and view-through duration, while higher image quality (produced by human-human teams) improved cost-per-click rates. We found three mechanisms explained these effects. First, human-AI collaboration was significantly more task-oriented, with 25% more task-oriented messages and 18% fewer interpersonal messages. Second, human-AI collaboration displayed more delegation, as participants delegated 17% more work to AI agents than to human partners and performed 62% fewer direct text edits when working with AI. Third, recognition that the collaborator was an AI moderated these effects as participants who correctly identified they were working with AI were more task-oriented and more likely to delegate work. These mechanisms, in turn, explained performance as task-oriented communication improved ad quality, specifically when working with AI, while interpersonal communication reduced ad quality; delegation improved text quality but had no effect on image quality (consistent with the jagged frontier) and was positively associated with diversity collapse, creating homogeneous outputs of higher average quality. The results suggest AI agents drive changes in productivity, performance, and output diversity by reshaping teamwork.
1 Introduction
Artificial intelligence (AI) tools have garnered attention for their potential to improve productivity and performance (Eloundou et al., 2024, Bick et al., 2024). For example, large language models (LLMs) decreased the average time taken for mid-level professional writing tasks by 40% and increased output quality by 18% (Noy and Zhang, 2023). For job seekers, AI assistance with resumes increased job hiring by an average of 8% (Wiles et al., 2023), and for customer support workers, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023). Moreover, productivity gains were greater for lower-skilled workers (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Choi and Schwarcz, 2023) and varied across task domains (Dell’Acqua et al., 2023). Evidence from an online labor market suggests there has already been a reduction in demand for freelance knowledge work with the advent of generative pre-trained transformer (GPT) models (Hui et al., 2023). In a study reviewing over 106 papers, Vaccaro et al. (2024) showed that human-AI groups outperformed humans alone in 85% of the studies.
While studies like Liu et al. (2023) explore levels of proactivity in AI agents, they typically do so without randomized controlled trials (RCTs) measuring productivity effects. The studies that use RCTs to estimate productivity effects tend to randomize access to LLM chatbots [e.g. Dell’Acqua et al. (2023) and Chen and Chan (2024)], which are not typically multimodal, do not include context, do not allow the chatbots to take independent actions or use APIs to call outside of the work environment, and do not provide a collaborative workspace where machines and humans can jointly manipulate output artifacts in real time. These innovations are meaningful because AI agents today possess all these features, yet the existing scientific literature studies none of them.
Furthermore, we currently lack fine-grained task level insight into how human-AI collaboration changes work processes and communication patterns and how these changes affect productivity and performance. The vast majority of existing research focuses on the productivity effects of GPT chatbots on individual workers (Noy and Zhang, 2023, Dell’Acqua et al., 2023) or how AI changes people’s perceptions, beliefs, and behaviors (Tey et al., 2024, Costello et al., 2024). However, it is unclear how these interactions evolve in real-time collaborations, especially in environments where AI agents can take autonomous actions, adapt dynamically to human input, and participate in tasks requiring creativity and coordination. Our lack of evidence on in vivo human-AI collaboration exists, in part, because off-the-shelf experimental platforms and studies do not provide collaborative workspaces where researchers can precisely record and measure the collaboration itself: e.g., transcripts of messages between machines and humans, logs of edits to output artifacts, and API (application programming interface) calls to outside agents or tools. In contrast, current AI applications, such as Notion AI and Cursor, already integrate AI agents into such collaborative workspaces and interfaces.
To address these gaps, we developed Pairit, a novel experimentation platform designed to study human-AI collaboration in real-world, extensible tasks. Pairit introduces several key innovations. It enables real-time collaboration between humans and AI agents, allowing participants to manipulate text, images, and workflows collaboratively in chat-enabled workspaces that mirror existing online AI collaboration work processes. The platform supports randomized pairings of humans and AI (i.e., human-human or human-AI teams) and allows for randomization of prompts and model fine-tuning. Critically, the AI can perform the complete set of equivalent actions that humans can perform in the collaborative workspace. In the marketing experiment analyzed in this study, these include sending chat messages, writing ad copy, editing ad copy, writing calls to action, editing calls to action, scrolling through images, editing images, selecting images, and generating new images using an external call to Dall-E 3. Moreover, Pairit captures every time-stamped keystroke, message, edit, swipe, scroll, selection, API call, and intermediate output, providing a rich datasets that allow for the detailed reconstruction of collaboration workflows. This represents a fundamental departure from the existing literature, which enables RCT-based evaluation of the outputs and productivity implications of human collaboration with LLM-based chatbots and co-pilots, but does not enable randomized experiments analyzing the task-level productivity and work process changes created in human collaborations with fully functioning, multimodal AI agents.
To study the work process, productivity, and performance implications of human collaboration with multi-modal AI agents, we conducted a large-scale randomized study of human-AI collaboration on advertising design and creation, a task requiring creativity, iteration, and precision. A total of 2,234 participants, representative of the U.S. population, were recruited through Prolific and randomly assigned to human-human or human-AI teams using the Pairit platform. Teams worked collaboratively to create marketing campaigns for a think tank’s year-end annual report, including generating and selecting ad images and writing ad copy and calls to action. This process was fully recorded, resulting in a dataset including 11,024 ads, 182,607 messages, 1,889,559 text edits, 62,119 image edits, and 10,074 AI-generated images, offering an unprecedented level of detail with which to understand work processes, communication, productivity and output quality.
Once the lab portion of the experiment was completed, we conducted a field experiment on the ads produced by human-human and human-AI teams. We obtained human and AI quality ratings of the ads, including the quality of the ad copy and images, and the human- and AI-evaluated likelihood of consumer engagement with the ads (measured by click-through rates). We then ran the ads in a real online display ad ecosystem, generating over 4.9 million impressions on X, and evaluated click-through rate, cost-per-click, view-through rate, and view-through duration metrics on the annual report using the platform’s ads API and DocSend’s view metrics, which allow us to record how much of the report consumers read, page by page, after clicking through on the ads.
Our analyses examine three performance effects documented in the AI literature but in an agentic collaboration context: productivity and performance (Noy and Zhang, 2023, Brynjolfsson et al., 2023), the jagged frontier of AI capabilities across tasks (Dell’Acqua et al., 2023, Gans, 2026), and diversity collapse, or the homogenization of AI-assisted outputs (Padmakumar and He, 2024, Chen and Chan, 2024, Hao et al., 2026). We found evidence of all three effects in our agentic setting: human-AI teams produced 50% more ads per worker, higher-quality text, and more homogeneous outputs, while human-human teams produced higher-quality images. Field tests revealed that higher image quality (from human-human teams) improved cost-per-click while higher text quality (from human-AI teams) improved click-through rates and view-through rates, with these offsetting quality effects explaining similar overall field experiment ad performance across team types.
Most importantly, we found that specific teamwork dynamics—task-orientation, delegation, and AI recognition—moderate these effects. While our treatment randomizes AI agents at the extensive margin (humans working with or without AI), the data we collected on communication patterns, delegation decisions, and AI recognition capture how people interact with, collaborate with, and delegate work to AI agents. In mining the rich collaboration data produced by Pairit, we found individuals in human-AI teams sent 62% more messages than those in human-human teams. Furthermore, human-AI teams sent 25% more content- and process-oriented messages, especially messages containing suggestions, instructions, prioritization, and planning. Conversely, human-human teams sent 18% more social and emotional messages, including messages that expressed rapport building, self-assessment, and concern. Task-orientation, when working with AI, was then associated with higher text quality, image quality and likelihood of clicking, while interpersonal communication was associated with lower text quality, image quality and likelihood of clicking. We also tracked work process changes: human-AI teams made 62% fewer direct edits to the copy and delegated 17% more work to their AI partners. Higher delegation was then associated with improved text quality but not image quality, and with reduced output diversity. To support the idea that AI reshapes teamwork, we found that participants who correctly recognized that they were working with AI were more task-oriented and more likely to delegate work to their AI partners, teamwork processes associated with higher performance. Together, these results strongly suggest that collaboration with AI agents reshapes work processes in specific ways and that these new teamwork dynamics moderate productivity gains and jagged frontier quality effects in human-AI collaboration.
Although prior studies have shown that AI tools can improve productivity and reduce task completion times (Noy and Zhang, 2023, Brynjolfsson et al., 2023), AI is often treated as a passive tool rather than as an active collaborator in prior research. As AI agents become integral to modern workflows, researchers are beginning to explore their role as work collaborators, rather than mere tools, emphasizing the importance of trust, transparency, and integration in human-AI partnerships (Makarius et al., 2020, Anthony et al., 2023, Collins et al., 2024). Our work contributes to this emerging literature by presenting the first task-level randomized experiment measuring the work process, productivity and performance implications of human-AI collaboration with fully functional, multimodal AI agents. The Pairit platform serves as the methodological infrastructure enabling this comparison, and the field experiment provides external validation of quality and performance differences. The core contribution is the empirical analysis of how human-AI collaboration reshapes task-orientation, delegation, communication, and recognition, as well as output quality, performance, and productivity. We hope these observations inform future research as AI agents become integral to the workplace.
2 Theory
2.1 Teamwork in Human-Human and Human-AI Teams
Effective teamwork requires more than individual competence and depends critically on how team members coordinate, communicate, and manage interpersonal relationships (Marks et al., 2001). These social dynamics generate collective intelligence that exceeds the sum on individual abilities (Woolley et al., 2010). A foundational distinction in this literature separates taskwork, the technical work itself, from teamwork, the processes that enable individuals to work together effectively (Salas et al., 1992). Teamwork encompasses three types of processes: transition processes where teams plan and set goals, action processes where they coordinate and monitor progress, and interpersonal processes where they manage conflict and build motivation (Marks et al., 2001). Teams cycle through these phases as they execute work.
Human teams invest substantial effort in all three. They build shared mental models through ongoing communication about goals, strategies, and progress, enabling them to anticipate each other’s actions (Mathieu et al., 2000). This shared understanding allows teams to negotiate who does what and when. Members also build trust, provide support, and resolve conflict to maintain cohesion, though this consumes time and cognitive resources.
The emergence of AI agents that can act autonomously within collaborative workspaces represents a fundamental shift from AI as a tool to AI as a teammate. Unlike chatbot-style interfaces where humans initiate each interaction, agents in our setting have full observability of the shared workspace and can read, write, and edit artifacts just as a human teammate would (Wang et al., 2024). This shift raises fundamental questions about how teamwork dynamics change when one team member is artificial (Seeber et al., 2020). Human-robot interaction offers relevant insights, though most of this work examines physical tasks with clearly delineated roles. Even so, consistent patterns emerge: in manufacturing, healthcare, and military contexts, people adapt their behavior based on perceived robot capabilities and reliability (Hancock et al., 2011), and trust calibration proves central to effective collaboration (Lee and See, 2004, Parasuraman and Manzey, 2010). Poorly calibrated trust leads to either over-reliance on flawed automation or underutilization of capable systems. These dynamics can enable human-AI combinations to outperform either alone, though the magnitude depends on task characteristics and collaboration design (Vaccaro et al., 2024, Collins et al., 2024). Knowledge work introduces additional complexity: AI agents can perform the same cognitive tasks as human teammates, creating ambiguity about optimal task allocation. At the organizational level, AI integration affects not just dyadic collaboration but broader structures of coordination and role design (Bankins et al., 2024).
The shift from human to AI teammates may alter the fundamental nature of collaboration. For example, human teammates bring idiosyncratic knowledge, experiences, and perspectives that create cognitive diversity (Hong and Page, 2004). This diversity allows teams to outperform homogeneous groups of higher-ability individuals, but only when social dynamics like turn-taking and mutual responsiveness allow diverse perspectives to surface (Woolley et al., 2010). These same social dynamics carry expectations of reciprocity, fairness, and recognition. AI teammates, by contrast, bring consistent availability and (within their capabilities) reliable execution, but without the diversity that emerges from different life experiences or the social fabric that binds human teams. We examine how these differences manifest in three process mechanisms: communication patterns, delegation behavior, and recognition of partner identity.
2.2 Task-Oriented vs Interpersonal Communication in Human-AI Teams
Human-human collaboration requires social maintenance, including building rapport, providing emotional support, and repairing conflicts when they arise (Marks et al., 2001, Hoegl and Gemuenden, 2001). These interpersonal processes are not merely overhead; they build the trust and psychological safety that enable effective coordination (Edmondson, 1999). Psychological safety, a shared belief that the team is safe for interpersonal risk taking, reduces the vulnerability that prevents team members from admitting errors or proposing untested ideas. Yet these processes also consume time and cognitive resources that could otherwise be devoted to task execution.
AI partners require none of this social maintenance. They do not take offense, need encouragement, or require relationship repair after disagreements. This absence of social needs may fundamentally shift communication patterns and change the nature of optimal communication for productivity and performance in human-AI teams. When AI agents can take actions on shared work, humans may shift from direct task execution to directing work through communication, adopting a more instrumental stance than they would with human partners (Tey et al., 2024). We therefore examine whether communication with AI teammates becomes more task-oriented and less interpersonal.
Communication patterns may also affect performance outcomes, not just differ across conditions. Team communication predicts performance, though the strength of this relationship depends on task characteristics and team composition (LePine et al., 2008). One reason is that interpersonal communication builds trust and enables implicit coordination, where members anticipate each other’s needs without explicit discussion (Rico et al., 2008). Yet these benefits come at a cost. Attention is a limited resource (Kahneman, 1973), and social maintenance consumes cognitive capacity that could otherwise support task execution. For tasks with clear objectives where implicit coordination matters less than explicit direction, redirecting attention from relationship management to task completion may improve productivity.
Because AI teammates do not require negotiation or consensus-building, task-oriented communication may prove more effective with AI than with humans. Conversely, interpersonal communication that builds rapport with human teammates may be counterproductive with AI partners, consuming attention without yielding the trust and coordination benefits it provides in human teams. We examine here whether task-oriented communication improves performance while interpersonal communication harms performance in human-AI teams.
2.3 Delegation in Human-AI Teams
When working with capable partners, individuals face choices about task allocation, deciding what to do themselves versus what to delegate. In human teams, social considerations shape these choices. Norms of fairness discourage assigning disproportionate workloads, concerns about overburdening partners limit requests, and expectations of reciprocity create implicit debts. Economic theory identifies a related constraint. Team production creates moral hazard because individual contributions are difficult to observe (Holmström, 1982). Delegation in human teams thus involves interpersonal negotiation and trust, not just task optimization.
AI teammates may alter these dynamics. Without relationships and social standing to maintain and protect, delegation to AI carries fewer interpersonal costs. Prior work shows people adjust their behavior with AI. For example, people respond differently to AI and human recommendations, exhibiting algorithm aversion in some contexts and appreciation in others (Dietvorst et al., 2015, Logg et al., 2019), and they reduce effort when AI predictions are confident (Agarwal et al., 2025, Fügener et al., 2022). Yet most of this work examines AI as advisor or predictor rather than as a collaborator taking actions on shared work. We examine delegation in settings where AI can take actions directly, editing text, selecting images, and communicating with human partners. Without the social costs of delegating to a peer, humans may delegate more readily to an AI agent. This shift from doing to delegating represents a fundamental change in how work is accomplished, from direct execution to supervision and direction (Morgeson et al., 2010, Seeber et al., 2020). We therefore examine whether humans delegate more work to AI collaborators than to human collaborators.
Beyond shifting how work is accomplished, delegation may affect output quality. Delegation effectiveness depends on the delegate’s competence (Leana, 1986). This principle extends to AI, which exhibits a jagged frontier of capabilities, excelling at some tasks while struggling with others (Dell’Acqua et al., 2023). We examine delegation effects across this frontier. Large language models are trained extensively on text and instruction-following, suggesting delegation may benefit text quality. Visual selection requires predicting audience response, a task where AI capabilities are less established, so delegation effects on image quality may be weaker or absent.
Delegation also has downstream consequences for output characteristics. More delegation means more of the final output reflects the partner’s contributions. When that partner is AI, delegation amplifies both the benefits of AI collaboration (Noy and Zhang, 2023, Brynjolfsson et al., 2023) and its costs, including homogeneity (Padmakumar and He, 2024) and limitations outside AI capabilities (Dell’Acqua et al., 2023). We therefore examine whether delegation is associated with improved output quality and whether these benefits vary with AI capability.
2.4 AI Partner Recognition
The task-oriented communication and delegation patterns that we argue improve human-AI performance assume that participants recognize their partner as AI. Yet people automatically apply social rules to computers (Nass and Moon, 2000, Reeves and Nass, 1996), defaulting to politeness, turn-taking, and concerns about fairness even when interacting with machines. This tendency may be counterproductive when it leads to social investment that yields no benefit with AI partners. Recognition may enable collaborators to override these defaults and match their behavior to the actual demands of the interaction.
Perceived partner identity affects trust, communication, and delegation patterns (Zhang et al., 2023, Glikson and Woolley, 2020), and appropriate reliance on automation depends on accurate assessment of a partner’s capabilities (Lee and See, 2004). Human-appropriate collaboration strategies include building consensus before acting, softening feedback to preserve relationships, and investing in rapport that enables future reciprocity. These strategies make sense with human partners who have feelings, memories, and participate in ongoing relationships, but they also consume effort without benefit when applied to AI. Participants who recognize their AI partner can instead adopt task-oriented communication (Section 2.2) and delegate more readily (Section 2.3), iterating quickly without social cost and focusing on task outcomes rather than relationship maintenance. Those who mistakenly believe their AI partner is human may achieve suboptimal outcomes not because AI collaboration is inherently limited, but because they apply strategies suited to a human partner rather than an AI partner.
Recognition may therefore improve performance by enabling behavioral adaptation. In our design, participants are not told whether their partner is human or AI until after the task, allowing us to measure the effects of the actual and perceived identity of the partner. We therefore examine whether recognition of an AI partner predicts adoption of the collaboration behaviors that we argue enhance human-AI performance, and whether these behaviors, in turn, affect output quality.
3 Methodology
Our study was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects. [^1] See osf.io/jfzha and osf.io/95dhu.
3.1 The Pairit platform
Once a participant enters the Pairit (Figure 1) platform, they are randomized into a queue for either the human-human or the human-AI condition. In the human-human condition, the participant collaborates with another human participant. In the human-AI condition, the participant collaborates with an AI agent. Pairit’s left panel is the task panel in which participants can participate in the task work process (in the context of this experiment to create, edit, and submit ads). The platform includes a carousel of seven pre-selected stock images provided by the think tank that the participants and the AI agent can choose from, and the participants (as well as the AI) can also access the Dall-E 3 application programming interface (API) to generate new images. The ad copy creation and editing panel includes a headline, primary text, and a description. All edits, including image selection, image generation, and ad copy changes, are synchronized in real-time across participants. The chat panel displays a “Typing…” indicator when the partner is composing a message. When participants submit an ad, the canvas clears and resets for the next submission. On the right is the chat panel in which participants can chat in real time with either another human participant or an AI agent. To the best of our knowledge, Pairit is the first platform to facilitate real-time collaboration between human-human or human-AI teams with team- and model-level randomization and real-time chat and editing on synchronized text and image interfaces in the context of multimodal human-AI collaboration workspaces.
Figure 1: The Pairit platform. On the left is the task panel, and on the right is the chat panel. In the human-human condition, chat messages and edits on the task panel (including text edits, image selections, and AI image generations) are synchronized in real-time. In the human-AI condition, the participant chats with an AI agent with full context of the user interface (UI; see Section 3.2), and the AI can edit text, select images, and generate AI images.
3.2 AI agent
Context.
The platform uses OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06), which processes text and images in a single model. To give the AI full context of the task and collaboration, each API call to gpt-4o is prompted with the information on the screen so it has the full context of the collaboration. In the prompt, we include the same text of the task given to participants, previous submissions, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, and general instructions. Moreover, a screenshot is taken of the image after every change and included as an input so the AI can observe the image and its evolution over time and edits. These details enable human collaboration with a fully functional, multimodal AI agent. The prompt template is shown in Appendix K.
Actions.
To ensure that the human-human and human-AI conditions are comparable, the AI agent can take the same actions a human participant except for submitting ads. The actions include sending messages, editing any element of the ad copy (including the headline, primary text, and description), selecting images, generating images using the Dall-E 3 interface, and waiting (i.e., not taking any action). The agent is prompted every 10 seconds whether to engage in action.
Chain-of-thought.
To ensure the agent is taking actions appropriately, we used chain-of-thought prompting (Wei et al., 2023). [^2] We used OpenAI’s structured output feature for chain-of-thought prompting. Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message and were developed and perfected through trial and error.
3.3 Procedure
AI randomization and queuing.
As soon as a participant was redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 2A). In the human-human condition, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey.
Figure 2: Overview of methods. (A) Participants are randomized into collaborating with another participant or an AI agent. (B) Participants collaborate with another participant or an AI agent to produce ads in a real-time collaborative workspace.
Pre-task survey.
After participants were paired with each other, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (Rammstedt and John, 2007).
Ad creation task.
Once the participants were paired and completed a pre-task survey, they entered the collaborative workspace in which they could message each other and edit and submit ads (Figure 2B). The edits were synchronized and messages were transmitted between the participants in real-time using websockets (i.e., tiptap.dev and pusher.com), as they are in commercial collaboration tools (e.g., Google Docs) and chat applications (e.g., Slack). The participants had 40 minutes to submit as many ads as they could produce. At the end of the 40 minutes, the participants were automatically redirected to the post-task survey.
Post-task survey.
After the completion of the ad creation task, the participants answered questions regarding their perception of AI, all on a 7-point Likert scale. These included two questions about their experience using AI (i.e., “I have used artificial intelligence (AI) chatbots before (e.g., ChatGPT, Bard).” and “I had a positive experience using AI chatbots.”), one question about their perception of their partner as an AI (i.e., “I believe my partner was an AI during the task.”) and a final question in which we revealed the identity of their partner (human or AI) and ask whether that changed the perception of the quality of the collaboration (i.e., “Your partner was [an AI assistant/a human]. Knowing this, to what extent has your perception of the quality of your collaboration changed?”). 85% of participants completed the post-task survey.
3.4 Participants
We pre-registered a sample of 2,500 participants from Prolific (www.prolific.com) based in the US, with representative stratification across gender and ethnicity. In total, 2,310 participants completed the task. Of the 2,500 participants who entered, 23 did not enter the matching queue (see Section 3.3 for details) and were removed from the study. A further 167 participants quit the study or were timed out before they were matched to a partner. In 83 human-human sessions (14% of HH teams), the matched partner did not complete the task due to dropout or timeout. Because these sessions do not represent collaborative work and submissions required confirmation by both participants, we exclude them from all analyses, yielding a final sample of 2,234 participants (1,258 human-AI, 976 human-human). The overall attrition rate was 8.9%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
3.5 Summary statistics and randomization checks
In total, the dataset includes 2,234 participants, 1,751 teams, 11,024 display ad submissions, 182,607 messages, 62,119 edits on images, 1,889,559 edits on ad copy, and 10,074 AI-generated images. To ensure that the randomization procedure successfully balanced covariates across experimental conditions, Table 1 compares key participant characteristics between the human-human and human-AI conditions. These include demographic variables (e.g. age, gender), participants’ work status, and the Big Five personality traits. No significant differences were detected across treatment conditions, indicating balanced randomization across participants.
Variables All Human-AI Human-Human t-statistic (SE) p-value
Individuals 2234 1258 976
Teams 1751 1258 493
Gender (% Male) 50.8% 50.4% 51.3% -0.438 (0.021) 0.661
Age 42 $\pm$ 14 43 $\pm$ 15 42 $\pm$ 14 1.296 (0.608) 0.195
Full-Time 47.7% 48.5% 46.7% 0.830 (0.021) 0.407
Part-Time 14.6% 13.4% 16.2% -1.883 (0.015) 0.060
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. For t-statistics, gender is coded as 1 for male and 0 for female, and employment status is 1 for the listed category. Balance checks for Big Five personality traits appear in Table A.
Table 1: Randomization checks
3.6 Incentives
To incentivize participants to create high-quality ads, we informed them that they were eligible for additional \$100 prizes based on the quantity and quality of the ads they submitted, as well as the performance of the ads in the marketing field experiment. This incentive structure was identical across human-human and human-AI conditions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances, but not if the ads are of low quality.” Ultimately, two participants were awarded \$100 each.
3.7 Message labeling
To label messages into categories, we prompt gpt-4o-mini-2024-07-18 for each message independently and ask for a category label. We enforce that the labels are from the set of pre-determined labels using OpenAI’s Structured Outputs API. Categories include content, process, social, emotional, feedback, and other. We group content and process messages as task-oriented communication, and social and emotional messages as interpersonal communication. See Appendix J for the prompts used in this analysis.
3.8 Evaluation of ad quality
Participants.
To obtain human ratings of ad quality, we recruited a separate sample of 1,995 participants from Prolific based in the US, with representative stratification across gender, age, and ethnicity. A total of 1,200 individuals entered the survey. Of these participants, 5 dropped out before submitting their surveys, resulting in a dropout rate of 0.42%. We built a custom platform for this survey, run on Google Cloud Platform’s App Engine. [^3] The code is available on GitHub. This evaluation was conducted from November 7 to 9, 2024, and took a median of 16.47 minutes to complete.
Ad samples.
To obtain ratings for all ads while avoiding survey fatigue, we created a random sample, with random order, of 40 ads per participant. To ensure that each ad received at least 3 ratings, we produced a set of 1,300 samples. As a participant entered our survey platform, we drew one sample from the set, without replacement, to provide to the participant, with each participant receiving a unique sample.
Ad mockups.
To obtain ratings of display ads that are as close as possible to display ads one would see on digital ad publishers, we first created mockups of each display ad. An ad mockup is a simulated representation of a display ad designed to closely resemble the appearance and functionality of an actual ad as it would appear on digital ad publishers’ platforms. We built a web application that populates the image, ad copy, shortened link, call-to-action, and other user interface items, including the Like, Comment, Share, and Close buttons, the profile picture, and the Sponsored tag. Screenshots were then programmatically taken of each mockup.
Survey items.
We used mockups of display ads. For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Figure F for an example of an item on the survey.
3.9 Field evaluation of ad performance
To assess real-world advertising outcomes, we ran ad campaigns on the social media platform X, pre-registered on osf.io/95dhu. We use the following as outcome variables: click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; as a fraction of document viewed), and view-through duration (VTD; in seconds). We tracked VTR and VTD by using a unique link on DocSend for each ad.
Due to the effects of divergent targeting in A-B testing (Braun and Schwartz, 2024), we follow recommendation 5 from Braun et al. (2024) and run a multi-ad study to test the causal effects of the ads on performance, conditional on online algorithms. We do not use a holdout in this study. Each ad has a unique DocSend link, so we expect the outcome in a hypothetical holdout group to be zero and attribute all document visits to the ads (Braun et al., 2024).
To run as many ads as possible while sampling evenly from human-human and human-AI teams and from the human predictions of ad quality, we implemented a stratified sampling approach. We obtained a sample of 2,000 ads from a total of 11,138 ads, sampling between one and two ads from each of the 1,751 teams. We first divided the ads into those created by human-human teams and human-AI teams and then further divided the ads into 10 strata according to the human predictions of the likelihood of user clicks, from low to high predicted click probability (see Section 3.8). We removed 8 ads that potentially violated moderation policies (due to the inclusion of e.g. violence, sexual, or drug content). We controlled for advertising spend in our analyses to account for auto-bidding. Consistent with other analyses, field analyses exclude 37 ads from human-human sessions where one partner did not complete the collaboration.
To prevent overlapping audience targeting, we assigned five unique ads to each of 400 campaigns, structuring them as 5-ad split tests within individual campaigns (noting the platform’s limit of five split tests per campaign). To further ensure no overlap across campaigns, we allocated a random set of 133 ZIP codes to each campaign, selecting ZIP codes with populations between 10,000 and 100,000.444The data is from https://data.census.gov/table/ACSDT5Y2020.B01003. We tested the robustness of this allocation across ZIP codes by conducting one-way ANOVA on population and income, yielding non-significant results: $F(399,53199)=0.954$, $p=0.734$ for population, and $F(399,53199)=0.973$, $p=0.636$ for income.
Ad impressions were delivered from January 21, 2025, to February 9, 2025. Due to platform limits in running separate ad campaigns, we ran 50 ad campaigns for two days at a time. We control for any potential temporal confounders with campaign random effects.
3.10 Model specifications
Human-AI collaboration effects.
Due to the experimental nature of our design, we measure the effects of human-AI collaboration using a standard regression model:
$Y_{i}=\delta\text{HAI}_{i}+\beta X_{i}+\epsilon_{i}\text{,}$ (1)
where $Y_{i}$ is the outcome, HAIi is 1 if the unit is in the human-AI condition and 0 otherwise, $X_{i}$ includes demographic controls (age, gender, Big Five personality traits), and $\epsilon_{i}$ is the error term. The unit of analysis $i$ varies by outcome: individual-level for messages, copy edits, image edits, AI image generations, submissions, message categories, and delegation rates; team-level for team submissions; and rating-level for ad quality scores. For team-level analyses, $X_{i}$ includes team-level controls (average age, gender composition, average personality traits).
Field evaluations.
To evaluate the results of our field experiment, we use the following mixed effects model:
$Y_{tc}=$ $\ \delta_{t}\text{HAI}_{tc}+\text{Text}_{tc}+\text{Image}_{tc}+\text{Click}_{tc}$ (2)
$+\text{Spend}_{tc}+u_{c}+\epsilon_{tc}\text{,}$
where $Y_{tc}$ represents the outcome (e.g., CPC, CTR, VTR, VTD) for ad $t$ in campaign $c$, Image, Text, and Click are human-rated quality measures from Section 3.8, $\text{Spend}_{tc}$ is the campaign spend, and $u_{c}\sim N(0,\sigma^{2}_{u})$ is the random intercept for campaign $c$.
Mediation analysis.
To trace the pathway from treatment to communication to production actions to quality outcomes, we estimate sequential regressions using Equation 1, substituting the relevant predictor and outcome at each step. For example, we regress production actions (i.e., delegation, submissions, manual edits) on communication patterns (i.e., task-oriented, interpersonal), and quality outcomes on delegation. This approach identifies which communication patterns are associated with production behaviors, and how delegation relates to quality outcomes.
Recognition of partner identity.
To test whether correct identification of partner type moderates treatment effects, we estimate:
$Y_{i}=$ $\beta_{0}+\beta_{1}\text{HAI}_{i}+\beta_{2}\text{Recognition}_{i}$ (3)
$\quad+\beta_{3}(\text{HAI}_{i}\times\text{Recognition}_{i})+\gamma X_{i}+\varepsilon_{i}\text{.}$
A significant $\beta_{3}$ indicates that the treatment effect differs depending on whether participants correctly identified their partner type.
To construct a measure of recognition, we use the partner perception question (“I believe my partner was an AI during the task”), measured on a 7-point Likert scale from strongly disagree to strongly agree. Using a median split (score $\geq$ 4 = “believed AI”), we define Recognition = 1 if participants correctly identified their partner type: believed AI in the H-AI treatment, or believed human in the H-H treatment.
Delegation.
We measure delegation as the fraction of text editing work performed by the partner rather than the focal participant. We defined delegation as one minus the participant’s character contribution divided by the session’s total character contribution, which, on average for human-human teams, is mechanically 50%. For human-AI teams, we detect AI-generated content by identifying large text jumps (simultaneous edits of more than 10 characters), which is how the AI produced the content compared to the human editing incrementally.
Output diversity.
To measure output diversity, we embed each submission’s ad copy (headline, primary text, and description) using OpenAI’s text-embedding-3-small model. We compute the centroid (mean embedding vector) for each treatment group and calculate the cosine distance from each submission to its own treatment’s centroid. Higher values indicate greater distance from the typical output in that condition, reflecting more diverse or atypical content. We aggregate to the user level by averaging across submissions and estimate treatment effects using Equation 1. This analysis is restricted to users who submitted at least one ad with text content (NN=2,034 of 2,234 participants).
4 Results
4.1 Productivity, Quality and Output Diversity
We begin by examining the primary outcomes of human-AI collaboration: productivity, ad quality, and output diversity. These outcomes reveal a consistent pattern: human-AI teams achieve higher individual productivity and better text quality, but at the cost of reduced image quality and less diverse outputs.
Collaborating with AI agents increases individual-level productivity.
Participants were incentivized to submit as many ads as possible within the time limit. Participants in human-AI teams submitted 50% more ads per worker, per unit time than their counterparts in human-human teams, with results consistent when demographic controls are added (Table 2). AI collaboration thus supports individual productivity by enabling participants to generate more output per worker compared to those in human-human teams.
Productivity Quality Output
Text Image Click Diversity
Intercept 4.601∗∗∗ 5.011∗∗∗ 4.666∗∗∗ 3.698∗∗∗ 0.303∗∗∗
(0.346) (0.038) (0.034) (0.035) (0.010)
Human-AI 1.954∗∗∗ 0.327∗∗∗ -0.133∗∗∗ -0.013 -0.094∗∗∗
(0.197) (0.024) (0.021) (0.021) (0.006)
Demographics Yes Yes Yes Yes Yes
Observations 2234 45803 45803 45803 2034
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Productivity is the number of submissions by participants. Text, Image, Click are human ratings (7-point scale) with standard errors clustered at the ad level. Output Diversity is cosine distance to treatment centroid (higher is more diverse). Robust standard errors account for heteroskedasticity. Full ratings model with covariates appears in Table G.
Table 2: Collaborating with AI agents improves productivity and text quality but lowers image quality (i.e., a jagged frontier), and decreases output diversity.
It is important to note that our study did not include a human-alone condition. Prior studies have consistently shown that AI tools enhance individual productivity compared to humans working alone across various tasks (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024, Wiles et al., 2023). In light of this evidence, our results suggest that the AI in human-AI teams acts as a partial substitute for an additional human collaborator. Without a human-alone condition, however, we cannot directly quantify AI’s incremental contribution beyond what an individual human might have achieved working alone. We explore this limitation further in the Discussion section.
With that caveat in mind, to test the robustness of our results to the alternative hypothesis that individual-level productivity differences are simply due to the submission of incomplete ads, we examined ad copy completion rates across conditions. As shown in Table E, participants in human-AI teams had consistently higher completion rates for ad copy elements compared to participants in human-human teams. Taken together, these findings suggest that human-AI teams are indeed more productive, in both the number of ads submitted and the ad completion rate.
Collaborating with AI agents improves text quality but lowers image quality.
In human evaluations of the quality of the ads, we found that human-AI teams produced higher quality ad text but lower quality ad images compared to human-human teams (Table 2). The estimated likelihood of clicking on the ad was indistinguishable between the two groups. The human ratings reveal that AI introduces important trade-offs in output quality. Specifically, collaborating with LLM-based agents improves text quality but reduces image quality. This pattern aligns with the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), where AI excels at some tasks (text generation) while underperforming at others (image quality prediction). These results were supported by AI evaluations of ad quality (see Table H) and may arise because LLMs are trained for next-word prediction rather than image quality prediction. [^5] Interestingly, in our AI evaluations of ad quality, we found that AI ratings were higher on the text and clicks for ads produced by human-AI teams and the same across the groups for image quality. In a way, it is unsurprising that the AI rated these ads’ text quality, image quality, and estimated click likelihood as equal to or better than those produced by human-human teams because these ads were created in collaboration with OpenAI’s gpt-4o. These results suggest that while LLMs enhance text-based outputs, their contributions to multimodal outputs like ad images may require complementary tools or fine-tuning for image-related tasks.
Collaborating with AI agents reduces output diversity.
Beyond productivity and quality, we examined whether collaboration with AI agents affects output diversity. Using cosine distance to treatment centroids as a measure of output diversity (Table 2), human-AI teams produce outputs that are closer to their condition centroid than human-human teams ($\beta=-0.094$, $p<0.001$), indicating a homogenization of teams outputs (Chen and Chan, 2024, Padmakumar and He, 2024). This diversity collapse suggests that while AI-generated content may be higher quality on average on certain dimensions, it lacks the variation that human creativity provides.
4.2 Ad Performance in the Field
Our findings on productivity and quality raise an important question about whether and how these trade-offs meaningfully impact actual performance. To evaluate the real-world performance of ads created by human-human and human-AI teams, we conducted a field experiment on the social media platform X. Our ad campaigns generated 4,932,373 impressions and 7,546 clicks over 20 days. This section examines how human-AI collaboration influences key advertising metrics, including cost-per-click (CPC), click-through rate (CTR), and view-through duration (VTD), reported in Table 3, and view-through rate (VTR) shown in Table I. The field study extends the lab experiment findings and tests how the distinct productivity and quality profiles of human-AI and human-human teams translate to advertising outcomes in a real world setting. Broadly, we found that image quality reduces CPC while text quality increases CTR and VTD. Since human-AI teams produce higher text quality but lower image quality than human-human teams, these effects are offset, and human-AI ads perform similarly to human-human ads overall.
CPC (\$) CTR (%) VTD (log-sec)
Intercept 11.628∗∗∗ -0.028 0.000
(0.705) (0.013) (0.121)
Human-AI 0.122 -0.000 0.002
(0.213) (0.004) (0.038)
Image -0.290 0.002 -0.002
(0.132) (0.003) (0.023)
Text -0.149 0.007∗∗∗ 0.037
(0.116) (0.002) (0.019)
Click 0.209 -0.002 -0.008
(0.148) (0.003) (0.027)
Spend Yes Yes Yes
Campaign RE Yes Yes Yes
Observations 1822 1962 4997
  • Notes: p<<0.1, p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Spend is controlled to account for auto-bidding. Campaign RE represents campaign random effects. CPC excludes ads with zero clicks. Robust standard errors account for heteroskedasticity.
Table 3: In the field experiment, higher image quality reduces cost-per-click while higher text quality increases click-through rate and view-through duration.
Click measures
We examined CPC and CTR using regression models with campaign random effects to account for unobserved heterogeneity across the 400 campaigns, with results in Table 3. For CPC, measured in dollars, our analysis reveals no significant effect of collaboration type. [^6] Higher spend consistently lowers costs and suggests divergent targeting and optimization, likely because we had kept on auto-bidding in our campaigns. Thus, we consider the effects present in this section as conditional on ad algorithms, as per recommendation 5 from Braun et al. (2024). Ads with stronger image quality significantly reduced CPC by \$0.29 or 2.5%. For CTR, expressed as a percentage, there was no direct human-AI effect. Text quality increased CTR by a small but statistically significant 0.007%. Thus, quality trade-offs between collaboration types translate to distinct click performance advantages.
View measures
We assessed view-through rate (VTR) and view-through duration (VTD; in log-seconds), using similar regression models, with VTD in Table 3 and VTR in Table I. Our analysis of VTD followed a similar pattern as observed in the click metrics. While there was no direct human-AI effect, text quality was associated with longer viewing duration ($p$ = 0.051). On average, one point higher text quality ratings were associated with approximately 4% longer view-through duration. Our analysis of VTR showed no significant effects for collaboration type or quality. These findings demonstrate that ads with higher text quality, which are more often produced by human-AI teams, are viewed for longer periods of time.
4.3 Teamwork and Collaboration Dynamics
Having established the outcomes of human-AI collaboration, we now examine differences in the collaboration process itself. Teams collaborate through two primary modes: communication and task execution. We analyze both to understand the shifts in teamwork dynamics underlying the productivity, quality and performance outcomes.
Collaborating with AI agents increases overall communication.
To test how human-AI collaboration affects communication, we measured the number of messages sent by participants in both human-human and human-AI groups. Participants worked collaboratively with and sent messages to their partners in real-time, for both human-human and human-AI groups. In the human-human group, participants messaged their human partners; in the human-AI group, participants messaged the AI. The Pairit platform logs a timestamped record of each message sent by a participant or AI. Participants in human-AI teams sent 62% more messages than those in human-human teams, with consistent results when demographic covariates were included (Table 4). These results indicate that collaborating with an AI partner encourages more frequent communication overall.
Communication Task Execution
Count Task-Oriented Interpersonal Copy Edits Delegation
Intercept 20.296∗∗∗ 0.362∗∗∗ 0.580∗∗∗ 1676.458∗∗∗ 0.478∗∗∗
(1.450) (0.014) (0.015) (57.646) (0.018)
Human-AI 12.580∗∗∗ 0.090∗∗∗ -0.107∗∗∗ -1032.406∗∗∗ 0.083∗∗∗
(0.843) (0.008) (0.008) (33.588) (0.010)
Demographics Yes Yes Yes Yes Yes
Observations 2234 2234 2234 2234 2233
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Count is the number of messages sent by a participant. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Copy edits are counts of text changes (i.e., text input events). Delegation is fraction of text work performed by partner. Robust standard errors account for heteroskedasticity.
Table 4: Collaborating with AI agents increases communication, with more task-oriented and less interpersonal messages. It also decreases direct edits to the ad copy while increasing the delegation to the AI agents.
Collaborating with AI agents shifts teams from interpersonal toward task-oriented communication.
In addition to the number of messages sent across conditions, we investigated whether the content of messages varied across conditions. We used gpt-4o-mini to label each message into categories, which we grouped into task-oriented and interpersonal communication (see Section 3.7). We found that human-AI teams sent 25% more task-oriented messages, while human-human teams sent 18% more interpersonal messages (Table 4). This shift indicates that collaboration with AI emphasizes task-oriented over interpersonal interaction, possibly because participants can focus more on the task without needing to navigate the social or emotional aspects typical of human collaboration.
Collaborating with AI agents shifts work from direct editing to delegation.
Participants in human-AI teams made 62% fewer direct copy edits compared to human-human teams (Table 4). However, participants in the human-AI condition still engaged in direct editing, indicating that while AI collaboration reduced the frequency of manual edits, it did not eliminate them entirely. Human-AI teams also delegated 58% of work to AI partners compared to 50% in human-human teams ($\beta=0.084$, $p<0.001$), a 17% increase. Unlike previous studies, such as Chen and Chan (2024), which artificially constrain participants to specific modalities when interacting with AI, our results show that in more unconstrained collaborative settings, where humans and agents can freely allocate effort to different tasks, participants utilized both direct involvement with the work product (ad editing) and interaction through messaging.
Taken together, these patterns in communication and task execution support a delegation workflow: humans direct AI partners rather than collaborating as peers. Human-AI teams communicate more frequently and with greater emphasis on task-oriented communication, while making fewer direct edits and delegating more work to their AI partners.
4.4 Mechanisms: Teamwork and Collaboration Dynamics Moderate Performance Outcomes
To address why collaborating with AI agents produces these outcomes, we conducted mediation analyses explained in Section 3.10. We found that participants who correctly recognized their AI partner engaged in more task-oriented and less interpersonal communication and delegate more work to their AI partner. We also found that directive communication was positively associated with more delegation and that both communication style and delegation predict output quality.
First, participants in the human-AI treatment group were highly accurate in identifying their partner as AI (94.5%), while human-human group participants were relatively less accurate (57.6%). When we tested whether this recognition shaped collaboration behavior (Equation 3) we found that correct identification predicts communication volume, style, and task execution (Table 5). Participants who correctly identified their partner as an AI agent sent more messages ($\beta_{3}=11.453$, $p<0.001$), with a higher proportion of task-oriented messages ($\beta_{3}=0.061$, $p<0.001$) and a lower proportion of interpersonal messages ($\beta_{3}=-0.059$, $p<0.01$). They also delegated more ($\beta_{3}=0.056$, $p<0.05$) and, despite making far fewer direct copy edits overall in human-AI teams ($\beta_{1}=-1155$, $p<0.001$), made slightly more edits than participants who did not correctly identify their partner ($\beta_{3}=150.6$, $p<0.05$), suggesting that recognition may prompt greater oversight of AI-generated content. Together, these findings indicate that awareness of AI capabilities prompts both a shift toward task-oriented, directive communication, more delegation, and a calibrated approach to task execution with greater oversight.
Communication Task Execution
Count Task-Oriented Interpersonal Copy Edits Delegation
Intercept 17.069∗∗∗ 0.371∗∗∗ 0.584∗∗∗ 1632.538∗∗∗ 0.483∗∗∗
(1.552) (0.017) (0.017) (69.997) (0.020)
Human-AI $\times$ 11.453∗∗∗ 0.061∗∗∗ -0.059∗∗ 150.632 0.056
   Recognition (1.676) (0.018) (0.019) (66.080) (0.023)
Human-AI 3.280 0.044∗∗ -0.066∗∗∗ -1155.053∗∗∗ 0.043
(1.317) (0.016) (0.016) (48.054) (0.020)
Recognition 3.904∗∗∗ -0.011 0.007 53.490 -0.012
(0.997) (0.013) (0.013) (61.577) (0.011)
Demographics Yes Yes Yes Yes Yes
Observations 2234 2228 2228 2234 2233
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Recognition indicates whether the participant correctly identified their partner type. Count is the number of messages sent by a participant. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Copy edits are counts of text changes (i.e., text input events). Delegation is fraction of text work performed by partner. Robust standard errors account for heteroskedasticity.
Table 5: Recognition of AI partner moderates communication and task execution.
Second, communication style and delegation predict productivity, ad quality and output diversity, but these effects vary across human-AI and human-human teams (Table 6, Panel A). In human-human teams, task-oriented communication is negatively associated with productivity ($\beta=-4.19$, $p<0.001$), text quality ($\beta=-0.416$, $p<0.05$), and image quality ($\beta=-0.267$, $p<0.05$). However, in human-AI teams, the reverse is true: task-oriented communication improves productivity ($\beta=4.69$, $p<0.001$) and ad quality when working with AI ($\beta=0.746$, $p<0.001$ for text; $\beta=0.506$, $p<0.01$ for image; $\beta=0.584$, $p<0.001$ for click). In contrast, while interpersonal communication is positively associated with productivity in human-human teams ($\beta=4.25$, $p<0.001$), it is negatively associated with productivity ($\beta=-4.80$, $p<0.001$) and ad quality in human-AI teams ($\beta=-0.685$, $p<0.001$ for text; $\beta=-0.581$, $p<0.001$ for image; $\beta=-0.599$, $p<0.001$ for click). These patterns suggest that interpersonal communication facilitates productivity in human-human teams, perhaps through rapport-building and coordination, whereas task-oriented, directive communication is more effective in human-AI teams.
Productivity Quality Output
Text Image Click Diversity
Panel A: Communication
Task-Oriented -4.187∗∗∗ -0.416 -0.267 -0.286 0.011
(0.847) (0.166) (0.126) (0.124) (0.027)
Task-Oriented $\times$ H-AI 4.693∗∗∗ 0.746∗∗∗ 0.506∗∗ 0.584∗∗∗ -0.035
(1.242) (0.206) (0.173) (0.166) (0.035)
Interpersonal 4.247∗∗∗ 0.269 0.211 0.194 0.014
(0.754) (0.156) (0.118) (0.119) (0.026)
Interpersonal $\times$ H-AI -4.799∗∗∗ -0.685∗∗∗ -0.581∗∗∗ -0.599∗∗∗ 0.030
(1.141) (0.195) (0.168) (0.161) (0.034)
Demographics Yes Yes Yes Yes Yes
Observations 2041 2041 2041 2041 2030
Panel B: Task Execution
Delegation -1.382 0.302∗∗∗ 0.132 0.178∗∗ -0.083∗∗∗
(0.586) (0.071) (0.069) (0.065) (0.012)
Demographics Yes Yes Yes Yes Yes
Observations 1222 1222 1222 1222 1222
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Panel A: both treatments; base coefficients are H-H. Panel B: H-AI only. Task-oriented is the fraction of messages labeled as content or process. Interpersonal is the fraction of messages labeled as emotional or social. Delegation is fraction of text work performed by partner. Productivity is the number of submissions by participants. Text, Image, Click are human ratings (7-point scale) with standard errors clustered at the ad level. Output Diversity is cosine distance to treatment centroid (higher is more diverse). Robust standard errors account for heteroskedasticity.
Table 6: Communication style and task execution predicts performance differently in human-AI vs. human-human teams.
In human-AI teams, delegation improves text ratings ($\beta=0.302$, $p<0.001$) and click ratings ($\beta=0.178$, $p<0.01$) but has no effect on image ratings ($\beta=0.132$, n.s.; Table 6, Panel B), consistent with the jagged frontier of AI capabilities (Dell’Acqua et al., 2023). Click ratings likely track text quality because clicking is partially driven by copy appeal, and delegation improves text without worsening image quality. Delegation also reduces output diversity ($\beta=-0.083$, $p<0.001$), explaining the diversity collapse observed in human-AI teams. Notably, delegation is associated with lower productivity ($\beta=-1.38$, $p<0.05$), perhaps because excessive delegators are less engaged overall. In summary, recognition shifts communication style and increases delegation, which in turn affects ad quality and diversity.
Illustrative examples.
To ground these quantitative findings, we extracted representative message exchanges. In H-AI teams, communication is directive: “Can you generate an image with a professional looking person at a desk?” followed by the AI responding with a generated image. Coordination appears as explicit task allocation: “Let’s divide this up. I’ll work on the headline and you handle the image.” In contrast, H-H teams show rapport-building: “Sorry for the delay! Had to grab coffee” … “No worries, take your time.” H-H teams also engage in collaborative negotiation: “What do you think about using blue instead?” … “Hmm, I was thinking green might work better because…” … “Good point, let’s try both and see.” These exchanges illustrate the directive versus collaborative dynamics that distinguish human-AI from human-human teams.
5 Discussion
5.1 Theoretical, managerial, and methodological implications
Theoretical implications.
This study extends teamwork theory to human-AI collaboration, demonstrating how LLM-based AI agents reshape work processes, communication dynamics, and task execution in collaborative settings (Schneider et al., 2021). Our findings connect to three effects documented in the AI literature: productivity gains (Noy and Zhang, 2023, Brynjolfsson et al., 2023), the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), and diversity collapse (Padmakumar and He, 2024, Chen and Chan, 2024). We find that human-AI collaboration induces a delegation workflow—task-oriented, directive communication and increased delegation to AI—which improves text quality (an AI strength) but not image quality (an AI weakness), with treatment effects attenuated by these jagged frontier effects. Delegation also reduces output diversity, as teams that delegate more produce outputs more similar to the typical human-AI output. This extends theoretical frameworks on human-AI collaboration design (Agarwal et al., 2025) and behavioral trust in AI teammates (Zhang et al., 2023) from decision-making contexts to agentic, multimodal creative work.
The shift from interpersonal to task-oriented communication aligns with work showing that humans apply social behaviors to machines (Reeves and Nass, 1996), but that these behaviors differ systematically from human-human interactions in that people interacting with AI become more demanding and instrumental (Tey et al., 2024). This shift reflects an optimal adaptation to AI partners that follow instructions reliably and require no social maintenance. While AI can streamline team processes (Wilson and Daugherty, 2018), our findings revealed trade-offs. In our experiment, use of AI enhanced ad text quality but reduced image quality. Our field experiment also revealed that these quality differences map to distinct performance outcomes. Higher text quality was associated with higher click-through rates and view-through duration, while higher image quality was associated with reductions in cost-per-click. This suggests the jagged frontier extends beyond output quality to ad performance in the field as AI capabilities create comparative advantages in different outcome metrics. Theories of human-AI collaboration must therefore account not only for task-specific competencies but also for how these competencies translate to heterogeneous downstream outcomes (Dell’Acqua et al., 2023).
Critically, we identify recognition, or participants’ awareness that their partner is an AI, as a key moderator of these dynamics. Participants who correctly identified their AI partner shifted toward more task-oriented communication and delegated more work, suggesting that awareness of AI capabilities triggers behavioral adaptation. This demonstrates that recognition is not merely perceptual but consequential: it reshapes collaboration behavior in ways that affect output quality. The asymmetry in recognition accuracy (94.5% in human-AI teams versus 57.6% in human-human teams) suggests that AI communication patterns are distinctive and detectable, which may itself shape how humans calibrate their collaborative strategies.
Moreover, our work challenges the traditional conceptualization of AI as a passive tool or assistant (Noy and Zhang, 2023, Dell’Acqua et al., 2023, Chen and Chan, 2024) by positioning LLM-based agents as active collaborators that shape team dynamics and, in turn, productivity and output quality (Anthony et al., 2023, Makarius et al., 2020). For example, the AI’s proficiency in generating ad copy influenced participants to focus on instructions rather than direct edits. Rather than merely augmenting human efforts, AI agents co-create outcomes, introducing new dynamics that affect team performance.
Our findings also contribute to the literature on team cognition and distributed expertise (Collins et al., 2024). In human-human teams, expertise is often shared through interpersonal exchanges, which foster trust but can slow decision-making. In contrast, human-AI teams in our study bypassed much of this social overhead. Interestingly, we found that interpersonal communication predicts higher output diversity, suggesting that social exchanges may facilitate the sharing of novel ideas between collaborators. This creates a tension: the efficiency gains from task-oriented communication come at the cost of creative variation. We show that traditional teamwork dimensions operate differently in human-AI teams: coordination becomes directive rather than mutual, conflict is avoided rather than resolved, and task allocation becomes delegation rather than negotiation. These patterns suggest that teamwork theory remains applicable to human-AI collaboration but requires adaptation for the asymmetry inherent in these teams.
Managerial implications.
Our work also offers actionable insights for managers and organizations seeking to integrate AI agents into collaborative workflows, particularly in creative and productivity-driven environments. Our findings demonstrate that AI agents significantly reduce social coordination costs, allowing team members to focus on task execution and content generation. Managers in industries like marketing, content creation, or consulting can leverage AI agents to streamline routine tasks, such as drafting campaign materials or reports, freeing human employees for strategic activities like audience analysis or brand positioning.
However, the study also highlights critical limitations that managers must address to fully harness AI’s potential in multimodal workflows. While LLM-based agents excelled in text generation, they underperformed in image-related tasks. This trade-off supports the idea that organizations should adopt a hybrid approach along the jagged frontier of AI capabilities (Dell’Acqua et al., 2023), allocating tasks within AI capabilities to agents and those outside AI capabilities, such as image selection, to humans or fine-tuning models to perform better on those tasks. Such an approach can mitigate the quality disparities observed in our study and is particularly relevant in domains where compelling visual elements drive downstream outcomes.
Furthermore, our results provide a framework for structuring human-AI teams to optimize productivity and performance. AI’s ability to handle bulk editing tasks allowed participants to engage in more content ideation and quality control. Managers can apply this division of labor to design efficient workflows, assigning AI agents to repetitive or data-intensive tasks while reserving human expertise for creative oversight and final approvals. To implement such workflows, organizations should invest in training programs that teach employees how to delegate tasks to AI effectively, fostering a collaborative culture that maximizes both technological and human contributions.
Finally, the study underscores the importance of aligning AI deployment with organizational goals and team dynamics. Our findings reveal that higher delegation to AI reduces output diversity, a concern for organizations that value creative variation. To mitigate diversity collapse, managers could design workflows that preserve human-driven ideation, such as human-only brainstorming sessions before AI-assisted execution. Although human-AI teams produce fewer total submissions than human-human teams, the 50% increase in individual productivity suggests that AI can serve as a partial substitute for additional personnel in certain contexts. Managers could use this insight to scale team output without proportionally increasing headcount, a cost-effective strategy for startups or resource-constrained firms. However, they must also be mindful of “collapsing onto the jagged frontier”: as teams rely more on AI, they may improve at tasks where AI excels (text generation) while atrophying at tasks where AI underperforms (image selection). Organizations should ensure that human skills in domains in which AI is weak are maintained through deliberate practice or task rotation.
Methodological implications.
This study advances the methodological landscape for studying human-AI collaboration through the development of the Pairit platform, a novel experimental framework that redefines how researchers investigate team dynamics in real-time, task-driven settings. Unlike traditional approaches that rely on aggregate performance metrics or isolated AI interactions (Noy and Zhang, 2023, Dell’Acqua et al., 2023), Pairit captures granular, time-stamped data on every keystroke, message, edit, and API call, enabling a detailed reconstruction of collaborative workflows. In our experiment, this capability allowed us to analyze 182,607 messages, 1,889,559 text edits, 62,119 image edits, and 10,074 AI-generated images, revealing nuanced patterns such as how AI-driven suggestions shifted human effort from direct text editing to instructional messaging. This methodological advance allowed us to analyze both aggregate and granular dynamics of human-AI teamwork and offers researchers new opportunities to discover insights regarding human-AI decision-making, task allocation, and interaction processes.
The randomization of team composition represents another significant methodological advance, moving beyond human-only versus human-AI comparisons to include human-human and human-AI pairings. Unlike prior research that often contrasts AI-augmented individuals with those working alone (Noy and Zhang, 2023, Vaccaro et al., 2024), our approach allowed us to evaluate the effects on collaboration itself, which is critical to understand given the proliferation of AI agents in workplaces. This randomization framework is also adaptable to other configurations, such as teams with multiple humans or multiple AI agents, enabling researchers to test how team composition influences outcomes in contexts like project management, educational group work, or distributed software development.
To fairly compare human-human and human-AI collaboration, another critical contribution is that AI agents are agentic. That is, AI agents on Pairit are capable of performing nearly all actions available to humans, including sending chat messages, editing text, selecting images, and generating new images via external APIs (e.g., DALL-E 3), with the sole exception of final ad submission in this study. This approach ensures direct comparability between human-AI and human-human conditions, a departure from prior studies where AI is often just a chatbot (Chen and Chan, 2024, Brynjolfsson et al., 2023). By enabling AI to act independently and actively, we analyze how AI’s active participation reshapes collaborative dynamics, which would not have been possible in experimental designs that treat AI as a passive tool.
Pairit’s multimodal capabilities further distinguish it as a methodological tool, supporting tasks that integrate text, images, and external API interactions. While our study focused on ad creation, Pairit’s architecture is extensible to diverse domains, such as software development (e.g., AI generating code while humans debug), scientific writing (e.g., AI drafting literature reviews), or multimedia content production (e.g., designing virtual reality environments). This versatility allows researchers to explore human-AI collaboration in tasks requiring varied cognitive and creative skills, addressing a gap in prior studies that often limit AI to text-based functions (Brynjolfsson et al., 2023).
These methodological contributions have broader implications for experimental research and organizational practice. Pairit’s ability to simulate realistic collaborative environments offers a testing ground for refining theoretical models of human-AI interaction and informing the design of AI-driven workflows. Ultimately, Pairit’s methodological innovations bridge the gap between controlled experiments and real-world applications and provide a robust, scalable framework for advancing our understanding of human-AI collaboration.
5.2 Limitations and future directions
Limitations.
While we conducted one of the first truly agent-randomized experiments in human-AI collaboration, combining laboratory and field tests, on a newly developed platform capable of recording fine-grained collaboration dynamics, our work is not without limitations. First, although we identify task-orientation, delegation, and recognition as mechanisms underlying human-AI collaboration dynamics, we cannot fully disentangle whether these patterns stem from the nature of AI communication or from human participants adapting their behavior upon recognizing their partner as AI. Our mediation analyses tracing the pathway from recognition to communication to delegation to quality are post hoc and correlational; while they are consistent with a causal pathway, alternative explanations remain possible. For instance, participants who are more comfortable with AI may both recognize it more readily and adopt more efficient collaboration strategies, with no causal relationship between recognition and behavior. Second, the limitations of current AI models, particularly vision models, present challenges for multimodal tasks. These models are optimized for next-word generation and specific visual tasks, such as identifying items in images, but are not designed for nuanced assessments like image quality prediction. This limitation likely contributed to the lower image quality observed in human-AI teams and underscores the need for AI systems tailored to specific creative and evaluative tasks. However, AI is a moving target and image selection and design competency likely will improve over time. As with any study of AI, results may drift as AI changes. Moreover, while our study provides a controlled experimental context, it may not fully capture the complexities of long-term collaboration with AI agents in real-world environments. Future research should explore how these dynamics evolve over extended periods and in more diverse task domains to validate and expand on our findings.
Finally, our study lacks a human alone condition, which prevents us from directly assessing how human-AI teams compare to individuals working independently. Without this baseline, we cannot precisely isolate AI’s marginal contribution to productivity and the coordination costs associated with collaboration. Nevertheless, our findings have practical implications for organizations looking to integrate AI agents into collaborative workflows, particularly in tasks requiring creativity and coordination. Our work lays a foundation for understanding human-AI collaboration in complex tasks, and future research can build on this by incorporating a human-alone condition to further disentangle marginal productivity effects.
Regarding generalizability, our findings should apply to tasks that share key features with ad creation: multi-modal creative tasks requiring both text and visual elements, where AI capabilities vary across modalities (the jagged frontier); iterative production tasks involving multiple rounds of generation, evaluation, and refinement; and settings with objective quality metrics. Our findings may be less applicable to high-stakes decision tasks where errors are costly and irreversible or tasks requiring deep domain expertise where AI capabilities are more limited. While our study captured within-session adaptation (humans learning AI capabilities through multiple ad iterations), longitudinal dynamics across repeated collaborations remain unexplored. The core mechanism we identify (task-oriented communication, higher delegation, and capability-dependent quality effects) should generalize broadly, but the direction of quality effects will depend on the specific task and AI capabilities.
Future directions.
Our work opens several promising avenues for future research. First, ours is one of the first randomized controlled trials to evaluate how AI agents affect work and communication processes. In our detailed, task-level analysis, we uncovered specific effects on communication, collaboration, and work processes, but these findings represent only a first step. Future work should provide more task-level evidence about how collaboration with AI agents changes communication, teamwork, work processes, and organization design, and how these effects drive productivity and performance across task and organizational contexts. Such research is essential to understanding the impact of AI on work and organizational performance. Second, the limitations of current vision models suggest a need for separate, specialized image generation models. Future studies could incorporate state-of-the-art visual models designed specifically for tasks like image quality prediction and generation, enabling a more nuanced understanding of AI contributions to multimodal creative tasks. Third, our findings are grounded in the context of ad design, but they invite exploration across a range of other collaborative domains. Important contexts include software development (e.g., coding), data analysis, collaborative writing, and financial accounting. Investigating Human-AI collaboration in these areas could reveal domain-specific dynamics and inform the broader design of AI systems tailored to various professional workflows. Finally, extending this research to longitudinal settings could provide insights into how Human-AI collaboration evolves over time. Long-term studies could examine the sustainability of productivity gains, the development of trust, and the potential for “learning effects” where humans adapt to working with AI agents or vice versa. These investigations would bridge the gap between controlled experimental settings and real-world applications, enhancing the generalizability of our findings. We intend to make Pairit freely available to academics advancing peer-reviewed research into human-AI collaboration. We believe this platform could support hundreds, if not thousands, of new experimental studies on the effects of AI on productivity, performance, and work processes.
5.3 Conclusion
Our work provides novel insights into how AI agents reshape work processes, productivity, and performance in collaborative settings. Our findings connect to three effects documented in the AI literature. First, human-AI teams achieved 50% higher productivity per worker, suggesting that AI agents can substitute for human collaborators. Second, we observed the jagged frontier of AI capabilities: text quality improved in human-AI teams while image quality declined, with these offsetting effects explaining similar field performance. Third, we find diversity collapse: teams that delegated more to AI produced outputs more similar to the typical human-AI output. We then found three mechanisms underlying these effects: task orientation (human-AI teams sent 25% more task-oriented messages and 18% fewer interpersonal messages), delegation (participants delegated 17% more work to AI and made 62% fewer direct edits), and recognition (participants who correctly identified their AI partner were more task-oriented and delegated more). Together, these mechanisms constitute a delegation workflow in which humans direct and delegate work to AI collaborators to unlock performance gains. These findings suggest that the productivity, quality, and output diversity effects of AI depend critically on how humans communicate with, collaborate with, and delegate work to AI agents.
6 Acknowledgments
We thank Dean Eckles and John Horton for their invaluable discussions. This work was supported by the Initiative on the Digital Economy at the Sloan School of Management at the Massachusetts Institute of Technology. The study was approved by the Massachusetts Institute of Technology institutional review board. The code and data are available upon request.
Appendix
Appendix A Big Five Personality Balance Checks
To complement the demographic balance checks reported in Table 1, we examined whether Big Five personality traits were balanced across experimental conditions. Table A shows no significant differences in openness, conscientiousness, extraversion, agreeableness, or neuroticism between human-human and human-AI conditions, confirming successful randomization.
Trait All Human-AI Human-Human t-statistic (SE) p-value
Openness 0.71 $\pm$ 0.20 0.71 $\pm$ 0.21 0.72 $\pm$ 0.19 -0.869 (0.009) 0.385
Conscientiousness 0.79 $\pm$ 0.18 0.78 $\pm$ 0.18 0.79 $\pm$ 0.17 -1.039 (0.008) 0.299
Extraversion 0.56 $\pm$ 0.21 0.55 $\pm$ 0.21 0.56 $\pm$ 0.21 -0.949 (0.009) 0.343
Agreeableness 0.70 $\pm$ 0.22 0.70 $\pm$ 0.21 0.69 $\pm$ 0.23 0.473 (0.010) 0.636
Neuroticism 0.48 $\pm$ 0.23 0.49 $\pm$ 0.24 0.47 $\pm$ 0.22 1.589 (0.010) 0.112
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Personality traits are normalized from a 7-point Likert scale.
Table A: Big Five personality trait balance checks
Appendix B Productivity
Table B reports the full regression results for productivity outcomes at both team and individual levels. At the team level, human-AI teams produced fewer total submissions, which is expected given that these teams have half as many human workers. At the individual level, however, participants in human-AI teams submitted 50% more ads than their counterparts in human-human teams.
Team Individual
Intercept 7.623∗∗∗ 3.850∗∗∗ 4.601∗∗∗
(0.277) (0.129) (0.346)
Human-AI -1.847∗∗∗ 1.925∗∗∗ 1.954∗∗∗
(0.315) (0.197) (0.197)
Demographics No No Yes
Observations 1751 2234 2234
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Robust standard errors account for heteroskedasticity.
Table B: Productivity of human-AI and human-human teams.
Appendix C Work Actions
Table C presents detailed results on work actions across experimental conditions. Participants in human-AI teams made substantially fewer direct copy edits but engaged in more image edits and AI image generations, suggesting a shift from direct text manipulation to delegation and visual content creation.
Copy Edits Image Edits AI Images
Intercept 1676.458∗∗∗ 28.397∗∗∗ 6.067∗∗∗
(57.646) (1.559) (0.390)
Human-AI -1032.406∗∗∗ 6.592∗∗∗ 0.328
(33.588) (0.931) (0.219)
Demographics Yes Yes Yes
Observations 2234 2234 2234
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Robust standard errors account for heteroskedasticity.
Table C: Collaborating with AI reduces copy edits but increases image edits and AI image generations at the individual level.
Appendix D Message Categories
Table D reports the distribution of message categories across experimental conditions. Messages were classified into content, process, social, emotional, feedback, and other categories using GPT-4o-mini. Human-AI teams sent significantly more content- and process-oriented messages, while human-human teams sent more social and emotional messages.
Content Process Emotional Social Feedback
Intercept 0.179∗∗∗ 0.187∗∗∗ 0.195∗∗∗ 0.391∗∗∗ 0.048∗∗∗
(0.012) (0.010) (0.010) (0.013) (0.005)
Human-AI 0.048∗∗∗ 0.040∗∗∗ -0.044∗∗∗ -0.067∗∗∗ 0.022∗∗∗
(0.006) (0.006) (0.005) (0.007) (0.003)
Demographics Yes Yes Yes Yes Yes
Observations 2228 2228 2228 2228 2228
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Robust standard errors account for heteroskedasticity.
Table D: Collaborating with AI increases content-, process-, and feedback-related messages while decreasing social and emotional messages.
Appendix E Ad Copy Completion
Table E examines whether productivity differences might be driven by submission of incomplete ads. Participants in human-AI teams had consistently higher completion rates for all ad copy elements (headline, primary text, and description), indicating that the productivity gains reflect genuine output rather than incomplete submissions.
Headline Primary Text Description
Intercept 0.808∗∗∗ 0.769∗∗∗ 0.719∗∗∗
(0.025) (0.026) (0.027)
Human-AI 0.150∗∗∗ 0.169∗∗∗ 0.161∗∗∗
(0.014) (0.014) (0.015)
Demographics Yes Yes Yes
Observations 2234 2234 2234
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Robust standard errors account for heteroskedasticity.
Table E: Individuals in Human-AI teams submit ads with more copy completed.
Appendix F Survey Interface
Figure F shows the user interface for the ad quality survey described in Section 3.8.
Figure F: The user interface for the ad quality survey.
Appendix G Human Ratings with Full Covariates
Table G presents the complete regression results for human quality ratings, including demographic controls and work process covariates.
Text Image Click
Intercept 5.011∗∗∗ 5.045∗∗∗ 4.666∗∗∗ 4.413∗∗∗ 3.698∗∗∗ 3.539∗∗∗
(0.038) (0.048) (0.034) (0.045) (0.035) (0.044)
Human-AI 0.327∗∗∗ 0.432∗∗∗ -0.133∗∗∗ -0.165∗∗∗ -0.013 0.027
(0.024) (0.052) (0.021) (0.044) (0.021) (0.044)
# Submissions -0.028∗∗∗ -0.012∗∗∗ -0.018∗∗∗
(0.002) (0.001) (0.001)
# AI Images Generated 0.008∗∗∗ 0.029∗∗∗ 0.023∗∗∗
(0.002) (0.002) (0.002)
# Human Messages 0.004∗∗∗ 0.002∗∗∗ 0.002∗∗∗
(0.000) (0.000) (0.000)
# Human Copy Edits 0.140∗∗∗ 0.042∗∗ 0.080∗∗∗
(0.017) (0.014) (0.015)
# Human Copy Edits $\times$ (H-AI) -0.032 -0.001 -0.019
(0.027) (0.027) (0.026)
# Human Image Edits -2.165 -1.054 -0.664
(0.893) (0.704) (0.722)
# Human Image Edits $\times$ (H-AI) 2.255 2.072 1.832
(1.004) (0.822) (0.827)
Demographics Yes Yes Yes Yes Yes Yes
Observations 45803 45803 45803 45803 45803 45803
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Standard errors are clustered at the ad level. Copy and image edits are per thousand.
Table G: Human ratings of ads with full covariates.
Appendix H AI Ratings with Full Covariates
AI ratings.
To obtain AI ratings of ad quality, we prompt OpenAI’s multimodal gpt-4o-mini-2024-07-18, which supports image input and structured outputs. [^7] See OpenAI’s documentation on vision and structured outputs. To make AI evaluations comparable to human evaluations, we asked the same questions of the AI as those given to human evaluators (see Section 3.8). Each item was evaluated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." See Appendix L for the full prompt.
Results.
Table H presents the complete regression results for AI quality ratings, using the same evaluation criteria as human raters. The treatment effects on text and image quality remain robust after controlling for participant characteristics and the number of edits made during the collaboration.
Text Image Click
Intercept 5.463∗∗∗ 5.396∗∗∗ 6.494∗∗∗ 6.369∗∗∗ 5.281∗∗∗ 5.253∗∗∗
(0.026) (0.038) (0.026) (0.034) (0.022) (0.032)
Human-AI 0.128∗∗∗ 0.120∗∗∗ -0.011 -0.114∗∗∗ 0.071∗∗∗ 0.040
(0.015) (0.031) (0.015) (0.030) (0.013) (0.027)
# Submissions -0.013∗∗∗ -0.010∗∗∗ -0.008∗∗∗
(0.001) (0.001) (0.001)
# AI Images Generated 0.024∗∗∗ 0.029∗∗∗ 0.020∗∗∗
(0.001) (0.001) (0.001)
# Human Messages 0.003∗∗∗ 0.002∗∗∗ 0.001∗∗∗
(0.000) (0.000) (0.000)
# Human Copy Edits -0.004 -0.005 -0.025
(0.011) (0.010) (0.010)
# Human Copy Edits $\times$ (H-AI) -0.038 0.004 -0.061∗∗∗
(0.018) (0.018) (0.016)
# Human Image Edits -1.460∗∗ -1.909∗∗∗ -1.259∗∗
(0.527) (0.446) (0.425)
# Human Image Edits $\times$ (H-AI) 0.342 2.381∗∗∗ 0.999
(0.581) (0.507) (0.477)
Demographics Yes Yes Yes Yes Yes Yes
Observations 11024 11024 11024 11024 11024 11024
  • Notes: p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Standard errors are clustered at the ad level. Copy and image edits are per thousand.
Table H: AI ratings of ads with full covariates.
Appendix I Field Experiment View Metrics
Table I reports the full regression results for view-through rate (VTR) and view-through duration (VTD) from the field experiment on X. While collaboration type showed no direct effect on view metrics, text quality was significantly associated with longer viewing duration, suggesting that the higher text quality produced by human-AI teams translates to greater audience engagement.
VTR VTD (log-sec)
Intercept 0.000 -0.000 -0.000 0.518∗∗∗ 0.000 0.000
(0.006) (0.012) (0.012) (0.053) (0.119) (0.121)
Human-AI 0.001 -0.000 0.012 0.002
(0.004) (0.004) (0.037) (0.038)
Click -0.000 -0.000 -0.008 -0.008
(0.003) (0.003) (0.027) (0.027)
Image -0.003 -0.003 -0.002 -0.002
(0.002) (0.002) (0.023) (0.023)
Text 0.002 0.002 0.037 0.037
(0.002) (0.002) (0.019) (0.019)
Spend 0.001∗∗∗ 0.001∗∗∗ 0.001∗∗∗ -0.001 0.004 0.004
(0.000) (0.000) (0.000) (0.002) (0.002) (0.002)
Campaign RE Yes Yes Yes Yes Yes Yes
Observations 4997 4997 4997 4997 4997 4997
  • Notes: p<<0.1, p<<0.05, ∗∗p<<0.01, ∗∗∗p<<0.001. Campaign RE represents campaign random effects.
Table I: Effects on view-through rate (VTR) and view-through duration (VTD) from the field study.
Appendix J Prompts for Message Labeling
The following is the Python code used to generate labels for each message independently:
from openai import OpenAI
from pydantic import BaseModel
from typing import Optional
from enum import Enum
client = OpenAI()
MODEL = "gpt-4o-mini-2024-07-18"
class CategoryLabel(str, Enum):
Content = "Content"
Process = "Process"
Social = "Social"
Emotional = "Emotional"
Feedback = "Feedback"
Other = "Other"
class Label(BaseModel):
category_label: CategoryLabel
def code(message):
system_message = ’’’
You are an expert at analyzing collaborative conversations.
For each message, label it with structured categories to reflect the conversation dynamics accurately.
Output the results in JSON format.
Label Categories:
- CategoryLabel:
- Content: The message shares information, facts, or deliverables directly related to the task.
- Process: The message addresses strategies or approaches to performing the task and real-time organizational or logistical details for the session.
- Social: The message builds rapport or contains social interactions not directly related to the task.
- Emotional: The message expresses emotions or attitudes related to the session or task.
- Feedback: The message provides constructive feedback or evaluative comments on the task.
user_message = f’’’
Label the message using the CategoryLabel options above.
<message>{message}</message>
response = client.beta.chat.completions.parse(
model=MODEL,
messages=[
{"role": "system", "content": system_message},
{"role": "user", "content": user_message},
response_format=Label,
return response.choices[0].message.content
Appendix K Prompts for the AI Agent
The following is the string used as the prompt for the AI agent. Strings inside curly brackets \${} are formatted with the relevant data.
<Definitions>
<Action/> - The action taken by you, the "Bot". These include Wait’, Chat’, EditText’, SelectImage’, and GenerateImage’. These are provided in the action history and includes the timestamp (t=) of each action.
<Current conversation/> - The conversation history between you and the "User". This includes the timestamp (t=) of each chat message. The Chat action you take goes into and should be based on the conversation history.
<Current copy/> - The current ad copy you are working on. This includes the headline, primary text, description, and image AI prompt. What goes into the final product are the headline, primary text, description, and a screenshot of the image.
<Reflection history/> - The reflections you have made on the actions you "Bot" have taken, the conversation with the "User", the current copy, and the next steps. These are provided in the reflection history and includes the timestamp (t=) of each reflection. These are not shown to the "User". They are only for your internal evaluation. Examine if the "User" is engaged, if the ad copy is improving, and if the task is being completed.
<Action history/> - The actions you "Bot" have taken. These are not by "User". These include the action type and the timestamp (t=) of each action. These are provided in the action history.
</Definitions>
<Submission history>
These submission cannot be altered.
\${submissions}
</Submission history>
<Your features>
\${features}
</Your features>
<Current task>
\${task}
</Current task>
<Current copy>
<Headline>
\${headline}
</Headline>
<Primary text>
\${primaryText}
</Primary text>
<Description>
\${description}
</Description>
<Image prompt>
\${imageAIPrompt}
</Image prompt>
</Current copy>
<Elapsed time in seconds>
\${elapsedTime}
</Elapsed time in seconds>
<Bot action history>
\${actionHistory}
</Bot action history>
<Reflection history>
\${reflections}
</Reflection history>
<Current conversation>
\${messageHistoryText}
</Current conversation>
<Instructions>
You are "Bot", working on a collaborative task with a human partner "User" via texting. Your focus is to complete the assignment. You and the human partner would get paid.
For chats, emulate the informality of a text conversation. You may use lower case or texting shorthands like "nah" or "lol". So keep your responses short when appropriate. Do not reveal you are an AI. Pretend you are a human.
Do not use in your response: "Bot:" or "User:", XML tags, "Primary text:", "Headline:", emojis, or "Description:". Do not use markdown.
You are "Bot". Do not generate the same chat messages. Do not repeat the same actions except for "Wait". Wait to give "User" the time to process. If "User" is silent, you can prompt them with a question or a suggestion.
Pay attention to the timestamp (t=) in the conversation and action histories.
When you Chat’, you should respond based on the conversation history.
When you EditText’, you should make edits to the current copy based on the task, the current conversation, and the current copy. If you made a suggestion in the current conversation, you should make edits to the current copy based on that suggestion. The Primary Text should be short, one sentence max. The Description can be slightly longer, but still concise.
When you SelectImage’, you should select an image based on the task, the current conversation, and the current copy. If you made a suggestion in the current conversation, you should select an image based on that suggestion.
When you GenerateImage’, you should generate an image based on the task, the current conversation, and the current copy. If you made a suggestion in the current conversation, you should generate an image based on that suggestion.
DO NOT TAKE ANY ACTION WITHOUT CONSULTING "USER". PROMPT "USER" FOR CONFIRMATION BEFORE EACH ACTION.
You can delegate the action to "User" by asking them to take the action.
Explain what you are planning to take action on before you do it. Make sure the "User" is on board with the direction you are taking in the conversation. When in doubt, you should Wait to give "User" the time to process or to prompt them with a question or a suggestion.
DO NOT REPEAT ACTIONS, NOT EVEN SIMILAR ACTIONS.
To engage user, chat with them. Ask questions. Make suggestions. Provide feedback. Make sure the user is engaged in the conversation. If the user is silent, prompt them with a question or a suggestion. If the user is not engaged, you should Wait to give the user time to process or to prompt them with a question or a suggestion. Prioritize user engagement over actions.
</Instructions>
Appendix L Prompts for AI Ratings
The following is the Python code used for AI ratings:
from openai import OpenAI
from pydantic import BaseModel
from typing import Optional
from enum import Enum
client = OpenAI()
MODEL = "gpt-4o-mini-2024-07-18"
class AdPerformanceEvaluation(BaseModel):
text: int
image: int
click: int
def rating(image_url, task):
system_message = f’’’
You are an expert marketing assistant trained to evaluate the effectiveness of advertisements based on their potential for engagement (e.g., clicks) and conversion (e.g., reading time on the report).
<task>{task}</task>
user_message = f’’’
Evaluate the display ad based on the following criteria, providing a score from 1 to 7 for each:
1. Text: The text is present, clear, relevant, and engaging. 1 is strongly disagree, 7 is strongly agree.
2. Image: The image is visually appealing. 1 is strongly disagree, 7 is strongly agree.
3. Click: I am likely to click on this ad. 1 is strongly disagree, 7 is strongly agree.
Just provide the ratings for each category with no additional commentary.
response = client.beta.chat.completions.parse(
model=MODEL,
messages=[
{"role": "system", "content": system_message},
{"role": "user", "content": [
{"type": "text", "text": user_message},
{"type": "image_url", "image_url": {"url": image_url}}
temperature=0.0,
response_format=AdPerformanceEvaluation,
return response.choices[0].message.content